# Large Language Models in Mental Health Care: a Scoping Review

### v1

@@ 3.2.1 Distribution of Model Architectures and Prompting Algorithms
A significant amount of existing work has been dedicated to directly prompting GPT-3.5 or GPT-4 for mental health applications without training. Some examples include depression detection,[76, 75] suicide detection,[68] cognitive distortion detection,[61] and relationship counseling.[56, 57] In these studies, LLMs function as intelligent chatbots, engaging with users to provide a range of mental health services, including analysis,[19, 66, 59] prediction [65] and support.[46, 74, 48]
To enhance their effectiveness, methods like few-shot prompting, which presents the LLMs with a small number of task demonstrations before requiring them to perform a task, and chain-of-thought (CoT) prompting,[83] which prompt the model to generate intermediate steps or paths of reasoning when dealing with problems, are employed. Drawing from the success of CoT in NLP, novel approaches such as chain-of-empathy prompting,[51] which incorporates the insights from psychotherapy (i.e., the therapists’ reasoning process) to prompt the LLMs to generate the cognitive reasoning of human emotion, and diagnosis-of-thought prompting,[61] which prompt the LLMs to make a decision using three diagnosis stages (i.e., subjectivity assessment, contrastive reasoning, and schema analysis), have been proposed to further improve LLM performance in the mental health domain.
Another stream of research focuses on the further training or fine-tuning of general LLMs using mental health-specific texts. This approach aims to inject mental health knowledge into existing base LLMs, leading to more relevant and accurate analyses and support. Notable examples include MentaLLaMA [59, 60] and Mental-LLM,[65] which fine-tuned the LLaMA-2 model and the Alpaca[79]/FLAN-T5[84] model, respectively, using social media data for enhanced mental health predictions. Similarly, ChatCounselor [47] leveraged the Psych8k dataset,[47] comprising real interactions between clients and psychologists, to fine-tune the LLaMA model. Additionally, ExTES-LLaMA [52] employed the LLaMA model, fine-tuned on emotional support dialogues.
@@ 3.2.2 Fine-Tuning Techniques
Acknowledging the high cost and extensive time required to train LLMs from scratch,[85, 86, 87, 88] existing studies have uniformly adopted fine-tuning existing models on mental health data. This approach allows the models to acquire expert domain knowledge and evolve into dedicated mental health-focused LLMs.
All the studies that involved fine-tuning adopted the instruction fine-tuning (IFT) technique. IFT is a type of fine-tuning where a model is fine-tuned for an instruction-following task that involves instructing a model to perform another task. In contrast, in classical types of supervised fine-tuning, the model does not have access to instructions but is directly tuned to perform a single downstream task. IFT allows domain knowledge to be injected into LLMs while improving LLM’s capability to follow human instructions. For example, giving a conversation between clients and psychologist, ChatCounselor [47] used the prompts: "Your task is to identify the patient and counsellor in the conversation. Summarize the conversation into only one round of conversation, one query or description by the patient and one feedback by the counsellor", to prompt GPT-4 [35] to generate instruction-input-output triples, in which the instruction could be "If you are a counsellor, please answer the questions based on the description of the patient." The input could be the query/question from the patient and output could be the feedback/answer from the model.
Table 2: Overview of Mental Health-Related Datasets. +Dataset curated or compiled in used in "Dataset and Benchmark" studies. "K" denotes a thousand. Information not provided is marked by N/P. New A stands for not applicable.
Name
Intended Task
Mental Condition
Data Source Sample size Constructed by Reviewer experience License
@@
Mental health service platforms 22,346 questions with 56,063 answers Human Trained laymen N/P
SMHD [90]
Mental health condition detection
@@
ADHD1, anxiety disorders, ASD [^2] Autism spectrum disorders, bipolar disorders, BPD2, depression, eating disorders, OCD3, PTSD4, schizophrenia, SAD5
Reddit 4,399K posts Weak supervision N/P $C$-BY-4.0
D4 [91]
Depression diagnosis-oriented dialogues
depression
@@
Human-machine dialogue 1,339 dialogues Human Experts Research only
Esconv [92]
Emotional support conversation
New A
@@
Crowdworker chatting simulation 1,053 dialogues (31,410 utterances) Human Trained laymen $C$-BY-NC-4.0
DialogueSafety [93]
Dialogue safety
New A
@@
Online counseling platforms 7,935 multi-turn dialogues Human Experts MIT
Dreaddit [94]
Stress detection
N/A
@@
Reddit 3,551 post segments Human Trained laymen Publicly Available
DepSeverity [95]
Depression severity prediction
Depression
@@
Reddit 3,551 posts Human N/P Publicly Available
SDCNL [96]
Suicide ideation detection
New A
@@
Reddit 1,895 posts Weak supervision N/P N/P
CSSRS-Suicide [97]
Suicide risk prediction
New A
@@
Reddit 500 users Human Experts CC-BY 4.0
Depression_Reddit [98]
Mental health condition detection
Depression
@@
Reddit 800 posts Weak supervision N/P Publicly available
CLPsych15 [99]
Mental health condition detection
Depression & PTSD
Reddit 1,746 users Human N/P Research only
Ji et al. [100]
Suicide ideation & mental disorder detection
Depression & PTSD
@@
Twitter 866 users & 59,212 posts Weak supervision N/P Research only
SAD [72]
Stress cause detection
New A
@@
SMS messages 6,847 SMS-like sentences Human Laymen $C$-BY-4.0
CAMS [101]
Depression & suicide cause detection
Depression
@@
Reddit 5,051 posts Human Trained students Publicly available
EMPATHETICDIALOGUES [102]
Empathetic dialogue generation
New A
[…]
MultiWD [103]
Wellness dimensions detection
New A
@@
Reddit 3,227 posts Human Trained students & experts Research only
IRF [104]
Interpersonal risk factors detection
New A
@@
Reddit 3,523 cases Human Trained students Research only
GoEmotions [105]
Emotion classification
New A
@@
Reddit 58,009 cases Human Laymen Apache 2.0
MHP Reddit[106]
Counseling
New A
Reddit 9,501 question-answer pairs Human N/P MIT
EPITOME [107]
Empathetic response generation
New A
@@
Reddit, TalkLife[108] 10K interactions on empathy Human Trained laymen Publicly available
Therapist Q&A [109]
Cognitive distortions detection
New A
@@
Kaggle 2,531 patient speech samples Human N/P N/P
ExTES [52]
Emotional support dialogue
New A
@@
ChatGPT 11,177 dialogues LLM generation Trained laymen Publicly available
BDI-Sen [110]
Mental health condition detection
Depression
@@
Reddit 4,973 sentences Human Experts Research only
PsySym [111]
Mental health condition detection
Depression, anxiety, ADHD, bipolar disorder, OCD, PTSD, eating disorder
Reddit 83,779 sentences Human Trained laymen & experts N/P
NVDRS[112]
Suicide behavior detection
New A
@@
Mortality & incident reports 1,462 reports Human Experts N/P
EATD-Corpus [113]
Mental health condition detection
Depression
@@
Interview transcripts 162 interview audio & transcripts Weak supervision N/P Publicly available
ANEW[114]
Lexicon sentiment detection
New A
@@
English vocabulary 1K words Human N/P N/P
Warriner et al.[115]
Lexicon sentiment detection
N/A
@@
English vocabulary 13,915 lemmas Human Trained laymen N/P
NRC-VAD [116]
Lexicon sentiment detection
N/A
@@
English vocabulary 20K words Human Laymen N/P
@@
Ma et al.[19]
User comment analysis
New A
Reddit 2,917 posts Human N/P N/P
Psych8k[47]
Mental health support
New A
Real-life counseling 260 interviews with 8,187 instruction pairs N/P N/P Publicly available
Yao et al. [49]
Counseling
Postpartum Mood & Anxiety Disorders
Helpline 7,014 conversations with 65,062 messages Weak supervision N/P N/P
[…]
Suicide risk prediction & cognitive distortion detection
N/A
Weibo 2,159 comments Human Experts Publicly available
Bhaumik et al.[69]
Suicide ideation detection & personalized psychoeducation
New A
Reddit 232K posts N/P N/P N/P
Kumar et al.[58]
Counseling
N/A
Crowdworker chatting simulation 945 survey responses Human Laymen N/P
@@ 3.3 Dataset Characteristics
Among the 34 studies we reviewed, 36 datasets were identified. These datasets contain diverse data applicable to a wide range of intended mental health care tasks. Most datasets are dedicated to detection or classification tasks. This category includes tasks such as mental health condition detection, mainly focusing on depression and post-traumatic stress disorder (PTSD), [95, 98, 99, 94, 111] cause of stress detection, [72] and interpersonal risk factor prediction.[103] Another major group focuses on text-generation tasks, such as simulating counseling sessions,[89, 106] responding to medical queries and generating empathetic dialogue.[102, 107] The remainder concerns more specialized applications, such as analysis of user discussions of emotion-support LLMs [117] and dialogue safety exploration.[93]
Sources of these datasets are imbalanced, with 20 datasets (55.6%) collected from social media platforms, such as Reddit, Twitter, and Weibo. Eleven datasets relied on more controlled venues, such as mental health counseling or interview dialogues simulated by clinicians [93] and collected from real-life scenarios.[89, 91, 47] suicide behavior detection corpus collected from mortality and incident reports,[112] and helpline conversations.[49] Less frequently seen sources include LLM-synthesized data, [52] existing sentiment lexicons,[114, 115, 116], and crowdworker simulated conversation.[92, 102, 58]
Dataset sizes and units vary with sources and annotation methods. Those drawn from social media and created via web crawling with the need for annotation typically have larger sizes. In contrast, those composed of expert content contain fewer samples. For example, SMHD [90] contains 4,399K posts collected from Reddit, while D4 [91] only contains 1,339 dialogues reviewed by experts. [118]
Most datasets were crafted by manual work with either collection or annotation. Six used weak supervision, which uses heuristics or helper models to collect or annotate data, and can thus achieve much larger scales. For example, SWMH [90] obtained annotations through subreddit titles and was able to include over 50K samples. One study [52] utilized LLM - ChatGPT in their case, to generate data, and had trained laymen to review them.
Most datasets involved human review of the content regardless of their construction methods. Eight datasets explicitly mentioned the involvement of experts in the review process. Fourteen datasets did not specify the experience level of the annotators. The remaining datasets were reviewed by laymen or trained laymen, including those who might be crowdworkers such as Amazon Mechanical Turk (MTurk) workers. Most papers relied on publicly available datasets. Six datasets [19, 47, 49, 67, 69] were self-constructed. Only one of them,[49] curated in a "Dataset and Benchmark" study, has been released for public use. Most of these datasets are released under licenses that restrict their use to non-commercial purposes. Details and implications of these restrictions and licensing agreements are provided in Appendix 7.
@@ 3.4 Validation measures and metrics
For mental health applicability, different variants of F1 scores are the most predominant metric adopted across studies.[66, 59, 60, 49, 51, 61, 75, 78, 76] Accuracy, as another fundamental metric, is also extensively employed.[65, 66, 59, 60, 61, 69] Recall (also known as sensitivity) [59, 49, 51, 78, 69] and precision [49, 51, 75, 62, 69] are also widely used, and frequently reported with F1 and accuracy. Diagnosis-focused studies, while also heavily relying on F1 and Accuracy, tend to employ additional metrics such as Area Under the Receiver Operating Characteristic (AUROC),[78, 69] and Specificity [78] for a holistic understanding of LLM diagnostic validity.
@@
In evaluating language proficiency, BLEU,[44, 119, 50, 75, 54] ROUGE,[59, 46, 50, 52, 75] Distinct-N,[44, 46, 52] and METEOR [52] are widely used to assess alignment with human-like language, diversity of expression, and overall quality of the generated text. Advanced metrics such as GPT3-Score,[59] BARTScore,[59] and BERT-Score [49, 50, 54] are tailored to evaluate semantic coherence and relevance of text in specific contexts, addressing the demand for nuanced understanding. Perplexity [46, 52] quantifies model predictability and text naturalness, while Extrema and Vector Extrema [52] analyze the diversity and distribution in the embedding space to reflect the model’s linguistic creativity and depth. The adoption of these traditional language evaluation metrics appears to be driven by a lack of automated, efficient, and easily understandable metrics designed for evaluating the quality of free-text generation of LLMs for mental health care. Human evaluation is thus frequently used across studies.
@@
Among the 34 articles reviewed, 19 employed a combination of human and automated evaluations, 5 exclusively used human evaluations, and 10 relied solely on automated methods. However, there is no single unified evaluation framework. A few studies [44, 45, 51, 52, 54] adopted or adapted published evaluation standards or attributes discussed in previous studies,[107, 120, 121, 122, 93] yet none of these frameworks are not widely adopted. Some of these studies and most other studies added or designed from scratch their evaluation frameworks and selected attributes from interest from their own judgment.[60, 46, 47, 48, 50, 52, 53, 54] A wide range of properties/attributes are proposed to be evaluated, and the frequently overlapped ones are empathy,[44, 46, 49, 51, 53, 54, 56, 57] relevance,[44, 59, 46, 49, 52, 53, 54, 56, 57] fluency,[44, 59, 46, 50, 52, 56, 57] understanding,[47, 52, 53, 56, 57] and helpfulness.[44, 46, 52, 54, 56, 57] Specifically, in intervention applications, studies often measure user engagement with interventions and technology adoption. The rate of intervention acceptance,[73, 64, 58] frequency of smartphone usage,[73] adherence to digital journaling protocols,[74] willingness to use the intervention again,[57, 64, 58] and intervention adherence rates [73, 74, 57, 64, 58] are observed to evaluate the effectiveness of LLMs in these studies.
@@
We noticed some attributes, despite sharing names, differ in definition across studies. For instance, ’Informativeness’ varies in context – in some studies,[44, 47] it relates to the richness of LLM responses, while in others,[52] it measures the extent of individual elaboration on emotional distress. Expert-led assessments [48, 53, 77, 63, 57] prominently involve direct analysis of model outputs or expert questionnaire scoring. Reliability metrics like Cohen’s Weighted Kappa,[44, 49] Krippendorff’s Alpha,[48] Fleiss’ Kappa,[59, 52] and inter-annotator agreement [54, 56] have been essential for validating study instruments. The study designs vary, with the number of reviewers ranging from as few as three [75] to as many as fifty.[52]
@@
Table 3: Explanation of Metrics Used to Evaluated LLM-Generated Content for Mental Health Care. Metrics are ordered alphabetically for easy indexing. Only model-oriented metrics are shown for simplicity.
@@
Accuracy [65, 66, 59, 60, 61, 69]
Proportion of correct predictions in classification tasks.
Mental Health
AUROC [78, 69]
Measures the model’s performance in distinguishing between classes.
Mental Health
BERT-Score [123]/BARTScore [124]/GPT3-Score[125] [59, 49, 50, 54]
Evaluate the quality of text generation against reference texts using respective models.
Language
BLEU-N[126] [44, 119, 50, 75, 54]
Measures the overlap of n-grams between the generated text and reference texts.
Language
Distinct-N [44, 46, 52]
Measures the diversity of language by calculating unique n-grams.
Language
F1 Score [66, 59, 60, 49, 51, 61, 75, 78, 76]
A combination of precision and recall into a single metric, balancing the two.
Mental Health
Extrema/Vector Extrema [52]
Analyze the diversity and distribution in the embedding space of the texts.
Language
METEOR[127] [52]
Aligns and compares similar words between the generated text and reference translations.
Language
Perplexity [46, 52]
Measures how well a probability distribution predicts a sample. Lower is better.
Language
Precision [49, 51, 75, 62, 69]
The proportion of positive identifications that were actually correct.
Mental Health
Recall (sensitivity) [59, 49, 51, 78, 69]
The proportion of actual positives that were identified correctly.
Mental Health
ROUGE-N/L [128] [59, 46, 50, 52, 75]
Measures the overlap of n-grams or longest common subsequence between texts.
Language
Specificity [78]
The proportion of true negatives that are correctly identified.
Mental Health
@@
Consistency [52, 46]
Ensuring the chatbot’s responses are aligned with its defined role without contradictions.
Language
Empathy [44, 46, 49, 51, 53, 54, 56, 57]
The degree to which the chatbot shows warmth, sympathy, and concern towards the user’s situation.
Mental Health
Engagement [54]
The chatbot’s ability to maintain prolonged and engaging conversations with the user.
Mental Health
Exploration [45, 51, 56]
Indicators of the user’s attempts to investigate deeper into the topics brought up in the conversation.
Mental Health
Guidance [47, 57]
The chatbot’s effectiveness in giving clear, actionable instructions and guidance.
Mental Health
Helpfulness [46, 52, 54, 56]
The extent to which the chatbot contributes to reducing the user’s emotional distress and improving mood.
Mental Health
Informativeness [44, 47]
The extent to which the chatbot provides useful and relevant suggestions for the user’s issues.
Mental Health
Interpretation [45, 51]
The user’s effort to restate or reinterpret the problems presented, as indicated by linguistic markers.
Mental Health
Professionalism [60, 50]
Evaluating the chatbot’s responses for their logical and psychological accuracy.
Mental Health
Reflectiveness [47, 56, 57]
The chatbot’s skill in rephrasing, mirroring user inputs, and demonstrating active listening.
Mental Health
Relevance [44, 59, 46, 49, 52, 53, 54, 56, 57]
The chatbot’s ability to generate content related to the user’s specific problems.
Mental Health
Safety [119, 57]
Adherence to established safety norms and guidelines in the chatbot’s interactions.
Mental Health
Understanding [47, 52, 53, 56, 57]
The chatbot’s level of accurately grasping and reflecting the user’s experiences and emotions.
Mental Health
Coherence [50, 52, 56]

### v3

@@ 3.2.1 Distribution of Model Architectures and Prompting Algorithms
A significant amount of existing work has been dedicated to directly prompting GPT-3.5 or GPT-4 for mental health applications without training. Some examples include depression detection[71, 70], suicide detection[63], cognitive distortion detection[74], and relationship counseling[52, 53]. In these studies, LLMs function as intelligent chatbots, engaging with users to provide a range of mental health services, including analysis[15, 61, 55], prediction [60] and support[42, 69, 44].
To enhance their effectiveness, methods like few-shot prompting, which presents the LLMs with a small number of task demonstrations before requiring them to perform a task, and chain-of-thought (CoT) prompting[81], which prompt the model to generate intermediate steps or paths of reasoning when dealing with problems, are employed. Drawing from the success of CoT in NLP, novel approaches such as chain-of-empathy prompting[47], which incorporates the insights from psychotherapy (i.e., the therapists’ reasoning process) to prompt the LLMs to generate the cognitive reasoning of human emotion, and diagnosis-of-thought prompting[74], which prompt the LLMs to make a decision using three diagnosis stages (i.e., subjectivity assessment, contrastive reasoning, and schema analysis), have been proposed to further improve LLM performance in the mental health domain.
Another stream of research focuses on the further training or fine-tuning of general LLMs using mental health specific texts. This approach aims to inject mental health knowledge into existing base LLMs, leading to more relevant and accurate analyses and support. Notable examples include MentaLLaMA [55, 56] and Mental-LLM[60], which fine-tuned the LLaMA-2 model and the Alpaca[75]/FLAN-T5[82] model, respectively, using social media data for enhanced mental health predictions. Similarly, ChatCounselor [43] leveraged the Psych8k dataset[43], comprising real interactions between clients and psychologists, to fine-tune the Vicuna[77] model. Additionally, Zheng et al. [48] employed the LLaMA model, fine-tuned on emotional support dialogues.
@@ 3.2.2 Fine-Tuning Techniques
Acknowledging the high cost and extensive time required to train LLMs from scratch[83, 84, 85, 86], existing studies have uniformly adopted fine-tuning existing models on mental health data. This approach allows the models to acquire expert domain knowledge and evolve into dedicated mental health-focused LLMs.
All the studies that involved fine-tuning adopted the instruction fine-tuning (IFT) technique. IFT is a type of fine-tuning where a model is fine-tuned for an instruction-following task that involves instructing a model to perform another task. In contrast, in classical types of supervised fine-tuning, the model does not have access to instructions but is directly tuned to perform a single downstream task. IFT allows domain knowledge to be injected into LLMs while improving LLM’s capability to follow human instructions. For example, giving a conversation between clients and psychologist, ChatCounselor [43] used the prompts: ”Your task is to identify the patient and counsellor in the conversation. Summarize the conversation into only one round of conversation, one query or description by the patient and one feedback by the counsellor”, to prompt GPT-4 [31] to generate instruction-input-output triples, in which the instruction could be ”If you are a counsellor, please answer the questions based on the description of the patient.” The input could be the query/question from the patient and output could be the feedback/answer from the model.
Table 2: Overview of Mental Health-Related Datasets. +Dataset curated or compiled in ”Dataset and Benchmark” studies. ”K” denotes a thousand. Information not provided is marked by N/P. New A stands for not applicable.
Name Intended Task Mental Condition Data Source Sample size Constructed by Reviewer experience License
@@
PsyQA [87] Counseling New A Mental health service platforms 22,346 questions with 56,063 answers Human Trained laymen N/P
@@
SMHD [88] Mental health condition detection ADHD1, anxiety disorders, ASD2, bipolar disorders, BPD3, depression, eating disorders, OCD4, PTSD5, schizophrenia, SAD6 Reddit 4,399K posts Weak supervision N/P CC-BY-4.0
@@
D4 [89] Depression diagnosis-oriented dialogues depression Human-machine dialogue 1,339 dialogues Human Experts Research only
@@
Esconv [90] Emotional support conversation New A Crowdworker chatting simulation 1,053 dialogues (31,410 utterances) Human Trained laymen CC-BY-NC-4.0
@@
DialogueSafety [91] Dialogue safety New A Online counseling platforms 7,935 multi-turn dialogues Human Experts MIT
@@
Dreaddit [92] Stress detection N/A Reddit 3,551 post segments Human Trained laymen Publicly Available
@@
DepSeverity [93] Depression severity prediction Depression Reddit 3,551 posts Human N/P Publicly Available
@@
Haque et al. [94] Suicide ideation detection New A Reddit 1,895 posts Weak supervision N/P N/P
@@
CSSRS-Suicide [95] Suicide risk prediction New A Reddit 500 users Human Experts CC-BY 4.0
@@
Depression_Reddit [96] Mental health condition detection Depression Reddit 800 posts Weak supervision N/P Publicly available
CLPsych15 [97] Mental health condition detection Depression & PTSD Twitter 1,746 users Human N/P Research only
@@
Ji et al. [98] Suicide ideation & mental disorder detection Depression & PTSD Twitter 866 users & 59,212 posts Weak supervision N/P Research only
@@
SAD [67] Stress cause detection New A SMS messages 6,847 SMS-like sentences Human Laymen CC-BY-4.0
@@
CAMS [99] Depression & suicide cause detection Depression Reddit 5,051 posts Human Trained students Publicly available
EMPATHETICDIALOGUES [100] Empathetic dialogue generation New A Crowdworker chatting simulation $\sim$25K multi-turn dialogues Human Laymen CC-BY-NC-4.0
@@
MultiWD [101] Wellness dimensions detection New A Reddit 3,227 posts Human Trained students & experts Research only
@@
IRF [102] Interpersonal risk factors detection New A Reddit 3,523 cases Human Trained students Research only
@@
GoEmotions [103] Emotion classification New A Reddit 58,009 cases Human Laymen Apache 2.0
MHP Reddit[104] Counseling New A Reddit 9,501 question-answer pairs Human N/P MIT
@@
EPITOME [105] Empathetic response generation New A Reddit, TalkLife[106] 10K interactions on empathy Human Trained laymen Publicly available
@@
Therapist Q&A [107] Cognitive distortions detection New A Kaggle 2,531 patient speech samples Human N/P N/P
@@
ExTES [48] Emotional support dialogue New A ChatGPT 11,177 dialogues LLM generation Trained laymen Publicly available
@@
BDI-Sen [108] Mental health condition detection Depression Reddit 4,973 sentences Human Experts Research only
PsySym [109] Mental health condition detection Depression, anxiety, ADHD, bipolar disorder, OCD, PTSD, eating disorder Reddit 83,779 sentences Human Trained laymen & experts N/P
@@
NVDRS[110] Suicide behavior detection New A Mortality & incident reports 1,462 reports Human Experts N/P
@@
EATD-Corpus [111] Mental health condition detection Depression Interview transcripts 162 interview audio & transcripts Weak supervision N/P Publicly available
@@
ANEW[112] Lexicon sentiment detection New A English vocabulary 1K words Human N/P N/P
@@
Warriner et al. [113] Lexicon sentiment detection N/A English vocabulary 13,915 lemmas Human Trained laymen N/P
@@
NRC-VAD [114] Lexicon sentiment detection N/A English vocabulary 20K words Human Laymen N/P
@@
Ma et al.[15] User comment analysis New A Reddit 2,917 posts Human N/P N/P
Psych8k[43] Mental health support New A Real-life counseling 260 interviews with 8,187 instruction pairs N/P N/P Publicly available
Yao et al. [45] Counseling Postpartum Mood & Anxiety Disorders Helpline 7,014 conversations with 65,062 messages Weak supervision N/P N/P
Qi et al. [62]+ Suicide risk prediction & cognitive distortion detection N/A Weibo 2,159 comments Human Experts Publicly available
Bhaumik et al.[64] Suicide ideation detection & personalized psychoeducation New A Reddit 232K posts N/P N/P N/P
Kumar et al.[54] Counseling N/A Crowdworker chatting simulation 945 survey responses Human Laymen N/P
@@ 3.3 Dataset Characteristics
Among the 34 studies we reviewed, 36 datasets were identified. These datasets contain diverse data applicable to a wide range of intended mental health care tasks. Most datasets are dedicated to detection or classification tasks. This category includes tasks such as mental health condition detection, mainly focusing on depression and post-traumatic stress disorder (PTSD) [93, 96, 97, 92, 109], cause of stress detection [67], and interpersonal risk factor prediction [101]. Another major group focuses on text-generation tasks, such as simulating counseling sessions[87, 104] responding to medical queries and generating empathetic dialogue[100, 105]. The remainder concerns more specialized applications, such as analysis of user discussions of emotion-support LLMs [115] and dialogue safety exploration[91].
Sources of these datasets are imbalanced, with 20 datasets (55.6%) collected from social media platforms, such as Reddit, Twitter, and Weibo. Eleven datasets relied on more controlled venues, such as mental health counseling or interview dialogues simulated by clinicians [91] and collected from real-life scenarios[87, 89, 43]. Suicide behavior detection corpus collected from mortality and incident reports[110], and helpline conversations[45]. Less frequently seen sources include LLM-synthesized data [48], existing sentiment lexicons[112, 113, 114], and crowdworker simulated conversation[90, 100, 54].
Dataset sizes and units vary with sources and annotation methods. Those drawn from social media and created via web crawling with the need for annotation typically have larger sizes. In contrast, those composed of expert content contain fewer samples. For example, SMHD [88] contains 4,399K posts collected from Reddit, while D4 [89] only contains 1,339 dialogues reviewed by experts [116].
Most datasets involved human review of the content regardless of their construction methods. Eight datasets explicitly mentioned the involvement of experts in the review process. Fourteen datasets did not specify the experience level of the annotators. The remaining datasets were reviewed by laymen or trained laymen, including those who might be crowdworkers such as Amazon Mechanical Turk (MTurk) workers. Most papers relied on publicly available datasets. Six datasets [15, 43, 45, 62, 64] were self-constructed. Only one of them[45], curated in a ”Dataset and Benchmark” study, has been released for public use. Most of these datasets are released under licenses that restrict their use to non-commercial purposes. Details and implications of these restrictions and licensing agreements are provided in Appendix B.
@@ 3.4 Validation measures and metrics
For mental health applicability, different variants of F1 scores are the most predominant metric adopted across studies [61, 55, 56, 45, 47, 74, 70, 73, 71]. Accuracy, as another fundamental metric, is also extensively employed [117, 61, 55, 56, 44, 74, 64]. Recall (also known as sensitivity) [60, 55, 45, 47, 73, 64] and precision [45, 47, 70, 57, 64] are also widely used, and frequently reported with F1 and accuracy. Diagnosis-focused studies, while also heavily relying on F1 and Accuracy, tend to employ additional metrics such as Area Under the Receiver Operating Characteristic (AUROC)[73, 64], and Specificity [73] for a holistic understanding of LLM diagnostic validity.
@@
In evaluating language proficiency, BLEU[40, 117, 55, 46, 48, 70, 50], ROUGE[55, 42, 46, 48, 70], Distinct-N[40, 117, 42, 48], and METEOR [48] are widely used to assess alignment with human-like language, diversity of expression, and overall quality of the generated text. Advanced metrics such as GPT3-Score[55], BARTScore[55], and BERT-Score [45, 46, 50] are tailored to evaluate semantic coherence and relevance of text in specific contexts, addressing the demand for nuanced understanding. Perplexity [42] quantifies model predictability and text naturalness, while Extrema and Vector Extrema [48] analyze the diversity and distribution in the embedding space to reflect the model’s linguistic creativity and depth. The adoption of these traditional language evaluation metrics appears to be driven by a lack of automated, efficient, and easily understandable metrics designed for evaluating the quality of free-text generation of LLMs for mental health care. Human evaluation is thus frequently used across studies.
@@
Among the 34 articles reviewed, 19 employed a combination of human and automated evaluations, 5 exclusively used human evaluations, and 10 relied solely on automated methods. However, there is no single unified evaluation framework. A few studies [40, 41, 47, 48, 50] adopted or adapted published evaluation standards or attributes discussed in previous studies[105, 118, 119, 120, 91], yet none of these frameworks are widely adopted. Some of these studies and most other studies added or designed from scratch their evaluation frameworks and selected attributes from interest from their own judgment[56, 42, 43, 44, 46, 48, 49, 50]. A wide range of properties/attributes are proposed to be evaluated, and the frequently overlapped ones are empathy[40, 45, 49, 50, 53], relevance[40, 117, 42, 43, 70], fluency[40, 117, 55, 42, 46, 48], understanding[48, 49], and helpfulness[42, 48, 50, 53]. Specifically, in intervention applications, studies often measure user engagement with interventions and technology adoption. The rate of intervention acceptance[68, 59, 54], frequency of smartphone usage[68], adherence to digital journaling protocols[69], willingness to use the intervention again[53, 59, 54], and intervention adherence rates [68, 69, 53, 59, 54] are observed to evaluate the effectiveness of LLMs in these studies.
@@
We noticed some attributes, despite sharing names, differ in definition across studies. For instance, ’Informativeness’ varies in context – in some studies[40, 43] ([40] did not directly use informativeness, so it is not included in table 3), it relates to the richness of LLM responses, while in others[48], it measures the extent of individual elaboration on emotional distress. Expert-led assessments [44, 49, 72, 58, 53] prominently involve direct analysis of model outputs or expert questionnaire scoring. Reliability metrics like Cohen’s Weighted Kappa[40, 45], Krippendorff’s Alpha[44], Fleiss’ Kappa[55, 48], and inter-annotator agreement [50, 52] have been essential for validating study instruments. The study designs vary, with the number of reviewers ranging from as few as three [70] to as many as fifty[48].
@@
Table 3: Automated Evaluation Metrics and Human Evaluation Metrics Used in More Than One Study. Metrics are ordered alphabetically for easy indexing. Only model-oriented metrics are shown for simplicity.
@@
Accuracy [117, 60, 61, 47, 71, 64] Proportion of correct predictions in classification tasks. Mental Health
AUROC [73, 64] Measures the model’s performance in distinguishing between classes. Mental Health
BERT-Score [121]/BARTScore [122]/GPT3-Score[123] [55, 56, 45, 46, 50] Evaluate the quality of text generation against reference texts using respective models. Language
BLEU-N[124] [40, 117, 55, 46, 48, 70, 50] Measures the overlap of n-grams between the generated text and reference texts. Language
Distinct-N [40, 42, 48] Measures the diversity of language by calculating unique n-grams. Language
F1 Score [61, 55, 56, 45, 62, 47, 74, 48, 70, 73, 63, 71] A combination of precision and recall into a single metric, balancing the two. Mental Health
Precision [45, 62, 47, 57, 64] The proportion of positive identifications that were actually correct. Mental Health
Recall (sensitivity) [55, 45, 62, 47, 73, 64] The proportion of actual positives that were identified correctly. Mental Health
ROUGE-N/L [125] [55, 42, 46, 48, 70, 50] Measures the overlap of n-grams or longest common subsequence between texts. Language
Specificity [50, 73] The proportion of true negatives that are correctly identified. Mental Health